# Fast Point Cloud to Mesh Reconstruction for Deformable Object Tracking

### v2

@@ I Introduction
The task of creating meshes from point clouds is a long-studied problem in the computer graphics field. Works in 3D reconstruction generate a mesh for one point cloud [17, 8, 22, 10], while 4D reconstruction works generate meshes for a sequence of deforming point clouds [9, 12, 3, 14, 19]. The main focus of both works is the accurate generation of refined meshes and hence time efficiency is compromised. Furthermore, most of the 4D reconstruction works are based on learning the canonical (or template) mesh, which limits them to training a separate model for each object category with the main focus on humans and animals. In 3D reconstruction, on one hand, implicit algorithms [17] which use marching cubes are accurate but time inefficient, and on the other hand, explicit models [8, 22, 10] that deform one or more initial template meshes are speed efficient with poor accuracy and topology. Overall, prior works in 3D reconstruction do not apply to the time dimension and hence do not allow for the tracking of the vertex positions of the deformed meshes. This proves to be limiting in robotics applications.
In our work, to find a balance between speed and accuracy, we assume a template mesh can be generated from the first point cloud of the deforming point cloud sequence using a SOTA method, e.g. [17] which operates in the implicit domain and generates meshes of the right topology with a speed of . The main focus of our work is generating deformed meshes for the remaining deformed point cloud frames by continuously deforming and tracking the template mesh. In this work, we assume that within a sequence of deforming point clouds of an object, the topology does not change and hence, the template mesh has the same topology as the deformed point cloud of any frame within the sequence. To predict the mesh coordinates of the deformed mesh, we use our model, as illustrated in Fig. 1, which consists of an encoder and decoder for extracting an encoding from the deformed point cloud and a Real-NVP [6] component which deforms the template mesh given the deformed encoding. The resulting pipeline has an inference speed of 0.017 seconds for a template mesh of 3000 vertices and a point cloud of 5000 points on an Omen Desktop(24-core, 3.2GHz Intel Core i9-12900K CPU, 64GB memory, NVIDIA GeForce 3090 GPU with 24GB memory).
@@ II Related Work
This category of work reconstructs a separate mesh for each frame of point cloud input, lacking consistent tracking of vertex positions, which can be crucial for robotic applications. These works fall into two categories. In the first category, the vertex positions of a template mesh are modified using a deformation network to best fit the input point cloud that needs to be reconstructed [8, 22, 10, 20, 7].
@@
To move around vertices, three different methods are widely used: Neural ordinary differential equations(NODE) [5], Real-NVP [6], and MLP. MLP is not homeomorphic and hence does not guarantee manifoldness by the proof in the supplementary material of ShapeFlow [10]. While normalizing flows using NODE [5] are homeomorphic and guarantee manifoldness, they come at a cost of time inefficiency due to the integrals in their constructions. One pass through the construction takes 0.04s on our hardware. In contrast, a normalizing flow using Real-NVP [6] guarantees manifoldness at a speed of 0.017s.
@@
Some papers such as 3DN [22], ShapeFlow [10], and Deformation-Aware 3D Model [20] use a retrieval network to find a template mesh similar to the input point cloud from a mesh template bank. The disadvantage of 3DN [22] is that the deforming is done using an MLP which does not guarantee the preservation of the manifoldness of the mesh. Furthermore, in contrast to our work, these works do not take into account the deformations of one single instance but rather go from one object instance to another object instance. Some other deforming papers such as NMF [7], Point2Mesh [8] and Neural Parts [16] start with one or multiple sphere meshes as a template. Point2Mesh [8] fits a sphere to one point cloud during inference(which is also its training stage) over three hours which is unfeasible for downstream robotic applications. NMF [7] makes use of flow networks consisting of neural ordinary differential equations to find the flow direction of the sphere template vertices. Therefore, their work, while preserving topology can not reconstruct meshes of a higher genus than zero. Neural Parts [16] deforms five template spheres using Real-NVP [6]. Starting with multiple template meshes gives the possibility of reconstructing meshes of more complex topology such as a chair with a backseat hole (genus of one). However, this method suffers from disconnectivity between the five resulting component meshes and needs prior knowledge of the number of initial spheres.
@@
The second category of 3D mesh reconstruction papers, models implicit functions such as signed distance functions or an indicator function and then uses marching cubes to reconstruct them. For instance, Shape As Points [17] uses marching cubes and deep Poisson reconstruction[11] and works at a rate of which is the fastest 3D mesh reconstruction in the state of the arts to the best of our knowledge. Additionally, it can reconstruct meshes of arbitrary topology with thin structures.
@@
Another category of papers combines both the implicit function learning and the flow of the vertices[18]. This medical paper[18] is used for the reconstruction of body organs and the point correspondence e.g. livers of different patients. However, due to the template learning of their pipeline, they have a separate pre-trained model for each organ of the human body. Additionally, their inference time is 180.10s which is not feasible for a robotics case of use.
@@
In summary, the fastest model for 3D mesh reconstruction to our knowledge is Shape As Points [17] with a rate of which can also generate the right mesh topology. However, this method still falls short in speed, particularly in the field of robotic manipulation, and lacks the essential capability of vertex tracking throughout the deformation sequence.
@@
In the Computer graphics domain, a lot of work has been done in the 4D mesh reconstruction field [9, 12, 3, 14, 19, 21]. This consists of firstly, a sequence of point clouds of an object deforming (human or animal in motion, articulated object opening/closing, non-rigid object twisting) being fed to the network. Then, the network predicts a mesh for every point cloud. In some of these papers, the correspondence between meshes (vertex tracking) is also predicted, and in some not [9]. For time-efficiency purposes, some of the works in the 4D field reconstruct only one single mesh of either a canonical state or the first state. This single mesh is referred to as the template mesh. Sequentially like Occupancy Flow [14] or in parallel like CaDeX [12]), the template vertices are mapped to new point coordinates for the remaining point cloud frames. The mesh of each frame is reconstructed using the same faces as the template mesh along with the newly mapped vertex coordinates. For instance, the Occupancy Flow [14] paper and the LPDC [19] paper, use an occupancy field and the multi-resolution isosurface extraction algorithm to reconstruct the mesh of the first point cloud frame and use it as a template mesh. Subsequently, Occupancy Flow [14] uses normalizing flows consisting of neural ordinary differential equations(LPDC [19] uses MLP) to move around the vertices of the template mesh. In CaDeX [12], the canonical state which is an implicit field is predicted in parallel to vertex tracking. Then, the marching cubes algorithm is used to reconstruct the canonical implicit field as the template mesh. Subsequently, the meshes of other point cloud frames are reconstructed with vertex tracking. This method also uses normalizing flows to move around the vertex coordinates. Its normalizing flow uses Real-NVP [6] consisting of coupling blocks which is more time efficient than NODE [5]. Although CaDeX [12] achieves outstanding results, they have an individual pre-trained model for every specific object category because of their template-learning-based pipeline. For instance, there is a separate pre-trained model for animals, humans, laptops, doors, etc. Similarly, LPDC [19] is limited to a pre-trained model for the motion of only humans. Furthermore, these works are not fast enough for the field of robotics: CaDeX [12] takes 0.365s for its rest frames on our machine and 0.68s according to their own recordings. Occupancy Flow [14] takes 0.212s for its rest frames on our machine. The rest frame speed does not take into account the template reconstruction that is done only in the beginning.
The mainstream datasets used in these works Occupancy Flow [14], CaDeX [12], LPDC [19] are deforming human identities from DFAUST [2], articulated objects from Shape2Motion [23], humans and animals from DeformingThings4D [13] and warped objects using the Occupancy flow[14] deformable data generation code.
@@ III-A Auto-encoder
During training, an encoder and decoder (auto-encoder model) are firstly pre-trained on the reconstruction of sampled point clouds from our training data using chamfer distance loss reconstruction as can be seen in Fig. 1(a). Henceforth, we refer to this loss as the chamfer distance reconstruction loss (). The encoder learns how to encode vital features of a deformed point cloud such that it can be decoded into a reconstructed point cloud. Subsequently, the Real-NVP is trained with the pre-trained encoder on the same training data. We do ablation studies in Sec. IV-B to investigate whether the auto-encoder needs to be frozen while we train the Real-NVP. The pre-training is done in order to ensure that the encoding contains the important features of the point cloud. The final employed auto-encoder model is from [1]. In order to find a suitable architecture, the auto-pointcloud architecture [1] is compared with FoldingNet architecture [25] in Sec. IV-B.
@@ III-B Real-NVP
Similarly to prior works [16, 12], we use Real-NVP in a conditional way. The conditional Real-NVP learns how to deform the template mesh vertices conditioned on the encoding of the deformed point cloud such that it fits the deformed point cloud. Real-NVP consists of several conditional coupling blocks. Each coupling block is a continuous bijective function which makes Real-NVP homeomorphic. Homeomorphisms are continuous bijective maps between two topological spaces that conserve topological properties. The homeomorphic property of the Real-NVP guarantees a stable deformation of the template mesh without causing any holes in the deformed mesh. The coupling block operation which is a conditional affine transformation is explained in Eq. 1. For each coupling block, one random dimension of the coordinates is chosen and is redefined based on all of the three dimensions of the coordinates and the auto-encoder encoding (which is marked as in Eq. 1) while the two unchosen dimensions of the coordinates remain unchanged. In Eq. 1, the randomly chosen coordinate dimension is . and are sequential one-by-one convolutions with and without nonlinearity respectively. The coupling block is applied several times and gradually moves around the template mesh coordinates. The deforming mesh results from the deformed coordinates and shares the same face topology as the template mesh. The learning is done using the chamfer distance loss between the final deformed mesh output from the Real-NVP and the groundtruth deformed mesh as illustrated in Fig. 1(a). Henceforth, we refer to this loss as the chamfer distance deformation loss().
@@ III-E Dataset
For the training and testing datasets, the simulation of object deformation is automated using the code from [14]. The code applies random displacement fields to a 3x3x3 grid using a gaussian random variable and then obtains a continuous displacement field by interpolating between the grid points using RBF [15]. The random warping field can then be applied as many times as possible. This consecutive application of the same warping field creates a deformation trajectory containing steps as can be seen in Fig. 3(a).
@@ IV-A Dataset Generation
Additionally, in order to be able to demonstrate the manipulation capacity, six watertight mesh instances were taken from the ycb[4] dataset and four different datasets as explained below were generated using the deformation code. The six watertight meshes include a pair of scissors, a hammer, an orange, a dice, a foam brick, and a bleach-cleanser bottle. Dataset B and D are briefly visualized in Fig. 3(b)(c).
@@
| Work | Time | Method | Problem |
Categories |
| Ours |
0.019s |
Real-NVP |
Template deformation |
6 |
| CaDeX[12] |
0.365s |
Real-NVP |
Subproblem:Template deformation |
1 |
| Occflow[14] |
0.212s |
NODE |
Subproblem:Template deformation |
2 |
| topology[18] |
185.27s |
NODE |
Subproblem:Template deformation |
1 |
| NMF[7] |
0.14s |
NODE |
3D reconstruction |
shapenet |
| Point2mesh[8] |
3h |
MLP |
3D reconstruction |
one instance |
@@ IV-B1 Auto-encoder
In a first stage, the possibility of using pre-trained encoders from other works was considered. These pre-trained encoders are trained on some of the Shapenet categories (8-10 different categories) where the training data does not contain any deformations. The purpose of this experiment was to see whether such pre-trained models are generalizable to deformations. The reuse of pre-trained encoders would save time and allow our work to focus solely on the training of the conditional Real-NVP. The works that were considered include PCN [26], GRNET [24], and FoldingNet [25] which are the SOTA of point cloud reconstruction on benchmarks. As can be seen in Fig. 4, qualitatively, none of the pre-trained encoders worked well on the deformed donut which was an unseen category during training. Additionally, PCN [26] and GRNET [24] did not do well on the deformed car which was a seen category. The trained FoldingNet also failed to capture the tires of the car. Thus, in the following experiments, the auto-encoder was retrained on datasets containing deformations. Furthermore, we tried different variations of these auto-encoders by changing the encoding size in order to preserve more features from the deformed point cloud.
During the training of the encoder, two different architectures were tested, the FoldingNet [25] and the Auto-pointcloud [1]. Additionally, the encoding size of these architectures was also studied. The results in terms of and are reported in Table II. The most indicative metric of the final quality is because it compares the final output mesh with the groundtruth mesh.
Current literature states that independent one-by-one convolutions maintain permutation invariance of point clouds but do not capture local features. Architectures such as FoldingNet overcome this by applying poolings on local neighborhoods using KNNs. However, in Table II, our experiments show a decrease in performance in terms of chamfer distance when using FoldingNet [25] in comparison to Auto-pointcloud [1].
Moreover, the FoldingNet [25] architecture and Auto-pointcloud [1] have better performance in terms of when the encoding size is smaller. This is unintuitive as one would expect the deformation to be of better quality if the encoding keeps more information about the deformed point cloud. Our results could be due to slight overfitting in the case of a small encoding size. However, we believe that a larger encoding size would enhance the performance considerably if the model were to be trained on many more object categories. When trained on a very diverse training set, the encoding would have to keep a lot of information in order to be generalizable to many object categories.
@@
| Exp. | Set | Auto-encoder Arch. | Embedding Size | End2end | ||
| Exp1 |
B |
Auto2018 |
256 |
0.0010 |
0.0011 |
no |
| Exp2 |
B |
Auto2018 |
1024 |
0.0010 |
0.0012 |
no |
| Exp3 |
B |
Folding-net |
256 |
0.0011 |
0.0016 |
no |
| Exp4 |
B |
Folding-net |
1024 |
0.0011 |
0.0017 |
no |
| Exp5 |
B |
Auto2018 |
1024 |
0.0012 |
yes |
| Exp6 |
B |
Folding-net |
1024 |
0.0013 |
yes |
| The encoder is not frozen while the Real-NVP is trained. | ||||||
| Exp. | Train set | Valid set | Auto-encoder Arch. | Embed Size | End2end | ||
| Exp.7 |
A |
B |
Auto2018 |
1024 |
0.0002 |
0.0002 |
yes |
| Exp.8 |
B |
B |
Auto2018 |
1024 |
0.0007 |
0.0003 |
yes |
| Exp.9 |
C |
C |
Auto2018 |
1024 |
0.0002 |
0.0004 |
yes |
| Exp.10 |
D |
C |
Auto2018 |
1024 |
0.0003 |
0.0005 |
yes |
| The encoder is not frozen while the Real-NVP is trained. | |||||||
@@
| Scissors |
0.0004 |
| Hammer |
0.0003 |
| Dice |
0.0016 |
| Bleach-cleanser |
0.0006 |
| Brick |
0.0015 |

### v3

# vision.bib Fast Point Cloud to Mesh Reconstruction for Deformable Object Tracking

@@ I Introduction
The task of creating meshes from point clouds is a long-studied problem in the computer graphics field. Works in 3D reconstruction generate a mesh for one point cloud [shapeaspoints, point2mesh, 3DN, shapeflow], while 4D reconstruction works generate meshes for a sequence of deforming point clouds [LCR4, cadex, NDG, occflow, lpdc]. The main focus of both works is the accurate generation of refined meshes and hence time efficiency is compromised. Furthermore, most of the 4D reconstruction works are based on learning the canonical (or template) mesh, which limits them to training a separate model for each object category with the main focus on humans and animals. In 3D reconstruction, on one hand, implicit algorithms [shapeaspoints] which use marching cubes are accurate but time inefficient, and on the other hand, explicit models [point2mesh, 3DN, shapeflow] that deform one or more initial template meshes are speed efficient with poor accuracy and topology. Overall, prior works in 3D reconstruction do not apply to the time dimension and hence do not allow for the tracking of the vertex positions of the deformed meshes. This proves to be limiting in robotics applications.
In our work, to find a balance between speed and accuracy, we assume a template mesh can be generated from the first point cloud of the deforming point cloud sequence using a SOTA method, e.g. [shapeaspoints] which operates in the implicit domain and generates meshes of the right topology with a speed of . The main focus of our work is generating deformed meshes for the remaining deformed point cloud frames by continuously deforming and tracking the template mesh. In this work, we assume that within a sequence of deforming point clouds of an object, the topology does not change and hence, the template mesh has the same topology as the deformed point cloud of any frame within the sequence. To predict the mesh coordinates of the deformed mesh, we use our model, as illustrated in Fig. 1, which consists of an encoder and decoder for extracting an encoding from the deformed point cloud and a Real-NVP [realnvp] component which deforms the template mesh given the deformed encoding. The resulting pipeline has an inference speed of 0.017 seconds for a template mesh of 3000 vertices and a point cloud of 5000 points on an Omen Desktop(24-core, 3.2GHz Intel Core i9-12900K CPU, 64GB memory, NVIDIA GeForce 3090 GPU with 24GB memory).
@@ II Related Work
This category of work reconstructs a separate mesh for each frame of point cloud input, lacking consistent tracking of vertex positions, which can be crucial for robotic applications. These works fall into two categories. In the first category, the vertex positions of a template mesh are modified using a deformation network to best fit the input point cloud that needs to be reconstructed [point2mesh, 3DN, shapeflow, deformAware, nmf].
@@
To move around vertices, three different methods are widely used: Neural ordinary differential equations(NODE) [neuralODE], Real-NVP [realnvp], and MLP. MLP is not homeomorphic and hence does not guarantee manifoldness by the proof in the supplementary material of ShapeFlow [shapeflow]. While normalizing flows using NODE [neuralODE] are homeomorphic and guarantee manifoldness, they come at a cost of time inefficiency due to the integrals in their constructions. One pass through the construction takes 0.04s on our hardware. In contrast, a normalizing flow using Real-NVP [realnvp] guarantees manifoldness at a speed of 0.017s.
@@
Some papers such as 3DN [3DN], ShapeFlow [shapeflow], and Deformation-Aware 3D Model [deformAware] use a retrieval network to find a template mesh similar to the input point cloud from a mesh template bank. The disadvantage of 3DN [3DN] is that the deforming is done using an MLP which does not guarantee the preservation of the manifoldness of the mesh. Furthermore, in contrast to our work, these works do not take into account the deformations of one single instance but rather go from one object instance to another object instance. Some other deforming papers such as NMF [nmf], Point2Mesh [point2mesh] and Neural Parts [neuralparts] start with one or multiple sphere meshes as a template. Point2Mesh [point2mesh] fits a sphere to one point cloud during inference(which is also its training stage) over three hours which is unfeasible for downstream robotic applications. NMF [nmf] makes use of flow networks consisting of neural ordinary differential equations to find the flow direction of the sphere template vertices. Therefore, their work, while preserving topology can not reconstruct meshes of a higher genus than zero. Neural Parts [neuralparts] deforms five template spheres using Real-NVP [realnvp]. Starting with multiple template meshes gives the possibility of reconstructing meshes of more complex topology such as a chair with a backseat hole (genus of one). However, this method suffers from disconnectivity between the five resulting component meshes and needs prior knowledge of the number of initial spheres.
@@
The second category of 3D mesh reconstruction papers, models implicit functions such as signed distance functions or an indicator function and then uses marching cubes to reconstruct them. For instance, Shape As Points [shapeaspoints] uses marching cubes and deep Poisson reconstruction[poissonsurfacerec] and works at a rate of which is the fastest 3D mesh reconstruction in the state of the arts to the best of our knowledge. Additionally, it can reconstruct meshes of arbitrary topology with thin structures.
@@
Another category of papers combines both the implicit function learning and the flow of the vertices[topologyPreservation]. This medical paper[topologyPreservation] is used for the reconstruction of body organs and the point correspondence e.g. livers of different patients. However, due to the template learning of their pipeline, they have a separate pre-trained model for each organ of the human body. Additionally, their inference time is 180.10s which is not feasible for a robotics case of use.
@@
In summary, the fastest model for 3D mesh reconstruction to our knowledge is Shape As Points [shapeaspoints] with a rate of which can also generate the right mesh topology. However, this method still falls short in speed, particularly in the field of robotic manipulation, and lacks the essential capability of vertex tracking throughout the deformation sequence.
@@
In the Computer graphics domain, a lot of work has been done in the 4D mesh reconstruction field [LCR4, cadex, NDG, occflow, lpdc, RFNet-4D]. This consists of firstly, a sequence of point clouds of an object deforming (human or animal in motion, articulated object opening/closing, non-rigid object twisting) being fed to the network. Then, the network predicts a mesh for every point cloud. In some of these papers, the correspondence between meshes (vertex tracking) is also predicted, and in some not [LCR4]. For time-efficiency purposes, some of the works in the 4D field reconstruct only one single mesh of either a canonical state or the first state. This single mesh is referred to as the template mesh. Sequentially like Occupancy Flow [occflow] or in parallel like CaDeX [cadex]), the template vertices are mapped to new point coordinates for the remaining point cloud frames. The mesh of each frame is reconstructed using the same faces as the template mesh along with the newly mapped vertex coordinates. For instance, the Occupancy Flow [occflow] paper and the LPDC [lpdc] paper, use an occupancy field and the multi-resolution isosurface extraction algorithm to reconstruct the mesh of the first point cloud frame and use it as a template mesh. Subsequently, Occupancy Flow [occflow] uses normalizing flows consisting of neural ordinary differential equations(LPDC [lpdc] uses MLP) to move around the vertices of the template mesh. In CaDeX [cadex], the canonical state which is an implicit field is predicted in parallel to vertex tracking. Then, the marching cubes algorithm is used to reconstruct the canonical implicit field as the template mesh. Subsequently, the meshes of other point cloud frames are reconstructed with vertex tracking. This method also uses normalizing flows to move around the vertex coordinates. Its normalizing flow uses Real-NVP [realnvp] consisting of coupling blocks which is more time efficient than NODE [neuralODE]. Although CaDeX [cadex] achieves outstanding results, they have an individual pre-trained model for every specific object category because of their template-learning-based pipeline. For instance, there is a separate pre-trained model for animals, humans, laptops, doors, etc. Similarly, LPDC [lpdc] is limited to a pre-trained model for the motion of only humans. Furthermore, these works are not fast enough for the field of robotics: CaDeX [cadex] takes 0.365s for its rest frames on our machine and 0.68s according to their own recordings. Occupancy Flow [occflow] takes 0.212s for its rest frames on our machine. The rest frame speed does not take into account the template reconstruction that is done only in the beginning.
The mainstream datasets used in these works Occupancy Flow [occflow], CaDeX [cadex], LPDC [lpdc] are deforming human identities from DFAUST [dfaust], articulated objects from Shape2Motion [shape2motion], humans and animals from DeformingThings4D [DeformingThings4D] and warped objects using the Occupancy flow[occflow] deformable data generation code.
@@ III-A Auto-encoder
During training, an encoder and decoder (auto-encoder model) are firstly pre-trained on the reconstruction of sampled point clouds from our training data using chamfer distance loss reconstruction as can be seen in Fig. 1(a). Henceforth, we refer to this loss as the chamfer distance reconstruction loss (). The encoder learns how to encode vital features of a deformed point cloud such that it can be decoded into a reconstructed point cloud. Subsequently, the Real-NVP is trained with the pre-trained encoder on the same training data. We do ablation studies in Sec. IV-B to investigate whether the auto-encoder needs to be frozen while we train the Real-NVP. The pre-training is done in order to ensure that the encoding contains the important features of the point cloud. The final employed auto-encoder model is from [2018auto]. In order to find a suitable architecture, the auto-pointcloud architecture [2018auto] is compared with FoldingNet architecture [foldingnet] in Sec. IV-B.
@@ III-B Real-NVP
Similarly to prior works [neuralparts, cadex], we use Real-NVP in a conditional way. The conditional Real-NVP learns how to deform the template mesh vertices conditioned on the encoding of the deformed point cloud such that it fits the deformed point cloud. Real-NVP consists of several conditional coupling blocks. Each coupling block is a continuous bijective function which makes Real-NVP homeomorphic. Homeomorphisms are continuous bijective maps between two topological spaces that conserve topological properties. The homeomorphic property of the Real-NVP guarantees a stable deformation of the template mesh without causing any holes in the deformed mesh. The coupling block operation which is a conditional affine transformation is explained in Eq. 1. For each coupling block, one random dimension of the coordinates is chosen and is redefined based on all of the three dimensions of the coordinates and the auto-encoder encoding (which is marked as in Eq. 1) while the two unchosen dimensions of the coordinates remain unchanged. In Eq. 1, the randomly chosen coordinate dimension is . and are sequential one-by-one convolutions with and without nonlinearity respectively. The coupling block is applied several times and gradually moves around the template mesh coordinates. The deforming mesh results from the deformed coordinates and shares the same face topology as the template mesh. The learning is done using the chamfer distance loss between the final deformed mesh output from the Real-NVP and the groundtruth deformed mesh as illustrated in Fig. 1(a). Henceforth, we refer to this loss as the chamfer distance deformation loss().
@@ III-E Dataset
For the training and testing datasets, the simulation of object deformation is automated using the code from [occflow]. The code applies random displacement fields to a 3x3x3 grid using a gaussian random variable and then obtains a continuous displacement field by interpolating between the grid points using RBF [rbf]. The random warping field can then be applied as many times as possible. This consecutive application of the same warping field creates a deformation trajectory containing steps as can be seen in Fig. 3(a).
@@ IV-A Dataset Generation
Additionally, in order to be able to demonstrate the manipulation capacity, six watertight mesh instances were taken from the ycb[ycb] dataset and four different datasets as explained below were generated using the deformation code. The six watertight meshes include a pair of scissors, a hammer, an orange, a dice, a foam brick, and a bleach-cleanser bottle. Dataset B and D are briefly visualized in Fig. 3(b)(c).
@@
| Work | Time | Method | Problem | Categories |
| Ours | 0.019s | Real-NVP | Template deformation | 6 |
| CaDeX[cadex] | 0.365s | Real-NVP | Subproblem:Template deformation | 1 |
| Occflow[occflow] | 0.212s | NODE | Subproblem:Template deformation | 2 |
| topology[topologyPreservation] | 185.27s | NODE | Subproblem:Template deformation | 1 |
| NMF[nmf] | 0.14s | NODE | 3D reconstruction | shapenet |
| Point2mesh[point2mesh] | 3h | MLP | 3D reconstruction | one instance |
@@ IV-B1 Auto-encoder
In a first stage, the possibility of using pre-trained encoders from other works was considered. These pre-trained encoders are trained on some of the Shapenet categories (8-10 different categories) where the training data does not contain any deformations. The purpose of this experiment was to see whether such pre-trained models are generalizable to deformations. The reuse of pre-trained encoders would save time and allow our work to focus solely on the training of the conditional Real-NVP. The works that were considered include PCN [yuan2018pcn], GRNET [xie2020grnet], and FoldingNet [foldingnet] which are the SOTA of point cloud reconstruction on benchmarks. As can be seen in Fig. 4, qualitatively, none of the pre-trained encoders worked well on the deformed donut which was an unseen category during training. Additionally, PCN [yuan2018pcn] and GRNET [xie2020grnet] did not do well on the deformed car which was a seen category. The trained FoldingNet also failed to capture the tires of the car. Thus, in the following experiments, the auto-encoder was retrained on datasets containing deformations. Furthermore, we tried different variations of these auto-encoders by changing the encoding size in order to preserve more features from the deformed point cloud.
During the training of the encoder, two different architectures were tested, the FoldingNet [foldingnet] and the Auto-pointcloud [2018auto]. Additionally, the encoding size of these architectures was also studied. The results in terms of and are reported in Table II. The most indicative metric of the final quality is because it compares the final output mesh with the groundtruth mesh.
Current literature states that independent one-by-one convolutions maintain permutation invariance of point clouds but do not capture local features. Architectures such as FoldingNet overcome this by applying poolings on local neighborhoods using KNNs. However, in Table II, our experiments show a decrease in performance in terms of chamfer distance when using FoldingNet [foldingnet] in comparison to Auto-pointcloud [2018auto].
Moreover, the FoldingNet [foldingnet] architecture and Auto-pointcloud [2018auto] have better performance in terms of when the encoding size is smaller. This is unintuitive as one would expect the deformation to be of better quality if the encoding keeps more information about the deformed point cloud. Our results could be due to slight overfitting in the case of a small encoding size. However, we believe that a larger encoding size would enhance the performance considerably if the model were to be trained on many more object categories. When trained on a very diverse training set, the encoding would have to keep a lot of information in order to be generalizable to many object categories.
@@
| Exp. | Set | Auto-encoder Arch. | Embedding Size | End2end∗ | ||
| Exp1 | B | Auto2018 | 256 | 0.0010 | 0.0011 | no |
| Exp2 | B | Auto2018 | 1024 | 0.0010 | 0.0012 | no |
| Exp3 | B | Folding-net | 256 | 0.0011 | 0.0016 | no |
| Exp4 | B | Folding-net | 1024 | 0.0011 | 0.0017 | no |
| Exp5 | B | Auto2018 | 1024 | - | 0.0012 | yes |
| Exp6 | B | Folding-net | 1024 | - | 0.0013 | yes |
| ∗The encoder is not frozen while the Real-NVP is trained. | ||||||
| Exp. | Train set | Valid set | Auto-encoder Arch. | Embed Size | End2end∗ | ||
| Exp.7 | A | B | Auto2018 | 1024 | 0.0002 | 0.0002 | yes |
| Exp.8 | B | B | Auto2018 | 1024 | 0.0007 | 0.0003 | yes |
| Exp.9 | C | C | Auto2018 | 1024 | 0.0002 | 0.0004 | yes |
| Exp.10 | D | C | Auto2018 | 1024 | 0.0003 | 0.0005 | yes |
| ∗The encoder is not frozen while the Real-NVP is trained. | |||||||
@@
| Scissors | 0.0004 |
| Hammer | 0.0003 |
| Dice | 0.0016 |
| Bleach-cleanser | 0.0006 |
| Brick | 0.0015 |